Rahul needs a loan and is speaking to several lending agencies about the interest rates they would charge and the terms they offer. He particularly likes his local bank because he is being offered a nominal rate of 8%. But the bank is compounding bimonth ly (every two months). What is the effective interest rate that Rahul would pay for the loan?

A

8. 271%

B

8.392%

C

8.585%

D

8.1 77%
Answer #1
Effective interest rate is given by the formula:
EIR = (1+r/m)^m-1, where r = the annual interest rate and m = the
number of compounding periods within a year, which is 6 in this
case.
EIR = (1+0.08/6)^6-1 = 8.271%
Answer: Option [A]